OpenStudy (anonymous):

square root of 56x^5

OpenStudy (anonymous):

Do you have any idea how to do this? Just so I know where you stand.

OpenStudy (anonymous):

23467.7 if x is not there if there is an X in the equation then I can't hel p you

OpenStudy (anonymous):

Well they way you do this is actually really simple because you can't completely square root either the x or 56 we have to find a number that will. so we break it down \[\sqrt{4\times14\times x ^{4}\times x}\] we can square root the 4 and the x^4, everything else is left under the square root sign. So the answer is \[2x ^{2}\sqrt{14x}\]